HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

INFRARED SPECTROSCOPY AND MOLECULAR DYNAMICS OF FLUOROMETHANE : ATMOSPHERIC INTEREST.

\underline{H. ZIADI}, GSMA, UMR CNRS 7331, University of Reims Champagne Ardenne, Moulin de la Housse B.P. 1039, F-51687 Reims Cedex 2, France, Laboratory of

Spectroscopy and Molecular Dynamics, National Superior School of Engineering of

Tunis, 5 Av Taha Hussein, 1008 Tunis, Tunisia; \underline{M. REY}, \underline{B. GROUIEZ} , \underline{M. ROTGER} , GSMA, UMR CNRS 7331, University of Reims Champagne Ardenne, Moulin de la Housse B.P. 1039, F-51687 Reims Cedex 2, France; \underline{A.V NIKITIN} , Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics, SB RAS, 634055 Tomsk, Russia; and \underline{H. AROUI}, Laboratory of Spectroscopy and Molecular Dynamics, National Superior School of Engineering of Tunis, 5 Av Taha

Hussein, 1008 Tunis, Tunisia

Recent advances in both experimental and theoretical rovibrational molecular spectroscopy permit to analyze and monitor the composition of atmospheric and astrophysical environments. The results obtained lead to a better understanding of the photochemical processes that govern the equilibrium of the atmospheric gases. This challenge was enhanced by the constant increase of the aptitude of optical remotesensing missions. Quantitative detections of these gases require accurate measurements of line positions, line strengths and broadening coefficients of molecular rovibrational transitions using high-resolution spectroscopy. Methyl fluoride CH_{3}F being a minor molecular constituent of the Venus atmosphere^{1}, is also present in the earth atmosphere at low concentration. The present work is dedicated to the 2ν_{6}, 2ν_{3} and ν_{3}+ν_{6} bands of CH_{3}F with the aim of generating a complete line list for databases (line positions, intensities, and self-broadening coefficients). The spectra were recorded at room temperature for seven pressures of CH_{3}F from 1.0172 to 14.956 mbar using the Bruker IFS125HR Fourier Transform spectrometer located at GSMA in Reims. This study builds upon our previous work on this molecule, which focused on the ν_{2} and ν_{5} bands and was published in the Journal of Quantitative Spectroscopy and Radiative Transfer in April 2022^{2}.

1K. Narahari Rao. and CW. Mathews, Molecular Spectroscopy: Modern Research,(1972)

2H. Ziadi, O. Ben Fathallah, I. Ben Chouikha, F. Kwabia Tchana, X. Landsheere and H. Aroui, J.

Quant. Spectrosc. Radiat. Trans., 286, 108218, (2022).

p-number: p004  Submitted on Mon, 22 May 2023 10:18:05 +0200